# Escaping the Subprime Trap in Algorithmic Lending

### v1

@@ 6 Conclusion
This is the posterior variance estimate from a Normal model with known mean and prior hyperparameters $\{1,{\sigma^{2}_{i}}^{(0)}\}$. ∎
@@
First, we suppose that Bank $L$ lends to group $W$, and derive a threshold rule for the prior variance $\sigma_{B}$ holding $\sigma_{W}$ fixed such that this rule determines whether or not Bank $L$ will lend to group $B$.
@@
We then need to check Bank $L$’s $VaR$ constraint. It is:
$Pr\left(\Pi_{j}^{(t)}|A_{BL}=1<\rho_{j}\right)=\Phi\left(\dfrac{\rho_{L}-(\mu_{B}+\mu_{W})}{\hat{\sigma}^{(t)}_{B}+\sigma_{W}}\right)$
@@
$Pr\left(\Pi_{j}^{(t)}|A_{BL}=1<\rho_{j}\right)\leq\alpha\>\Longleftrightarrow\Phi\left(\dfrac{\rho_{L}-(\mu_{B}+\mu_{W})}{\hat{\sigma}^{(t)}_{B}+\sigma_{W}}\right)\leq\alpha$
@@
$\hat{\sigma}^{(t)}_{B}\leq\dfrac{\rho_{L}-(\mu_{W}+\mu_{B}+\Phi^{-1}(\alpha)\sigma_{W})}{\Phi^{-1}(\alpha)}$
Where the right-hand side defines a variance threshold $\tilde{\sigma_{B}}$ such that, if group B’s prior variance is above this threshold, Bank $L$ will not lend to group $B$:
$\hat{\sigma}^{(t)}_{BL}\leq\tilde{\sigma}_{B}\equiv\dfrac{\rho_{L}-(\mu_{W}+\mu_{B}+\Phi^{-1}(\alpha)\sigma_{W})}{\Phi^{-1}(\alpha)}$
@@
$\sigma_{W}\leq\dfrac{\rho_{L}-\mu_{W}}{\Phi^{-1}(\alpha)}$
Since we know that, by Assumption 3, this condition is satisfied, Bank $L$ is willing to lend to $W$ unilaterally. Assumption 3 also tells us that $\hat{\sigma}^{(0)}_{B}$ is above this threshold, indicating that Bank $L$ will not lend to $B$ unilaterally.
Hence, Bank $L$ will lend to group $W$ unilaterally, and will lend to group $B$ in addition to group $W$ if and only if group $B$’s variance satisfies $\hat{\sigma}^{(0)}_{B}\leq\tilde{\theta}_{L}$. ∎
6.1 Proof of Lemma 3
@@
$Pr\left(\Pi_{j}^{(t)}|A_{BL}=1<\rho_{j}\right)=\Phi\left(\dfrac{\rho_{L}-[1+\nu](\mu_{B}+\mu_{W})}{\hat{\sigma}^{(t)}_{B}+\sigma_{W}}\right)$
@@
$\hat{\sigma}_{BH}^{(t)}\leq\tilde{\sigma}_{H}\equiv\dfrac{\rho_{L}-[(1+\nu)(\mu_{W}+\mu_{B})+\Phi^{-1}(\alpha)\sigma_{W}]}{\Phi^{-1}(\alpha)}$
6.2 Proof of Corollary 1
@@
[…]
@@
In each period, the Bank has belief, $\pi_{B}^{(t)}\sim\mathcal{N}\bigl{(}\mu,(\hat{\sigma}_{BL}^{(t)})^{2}\bigr{)}$, so that:
$\mathbb{P}\!\left(\Pi_{j}^{(t)}+s<\rho_{L}\right)=\Phi\!\left(\frac{\rho_{L}-(\mu_{W}+\mu_{B})-s}{\hat{\sigma}_{BL}^{(t)}+\sigma_{W}}\right).$
Thus, to guarantee $\Phi\!\left(\frac{\rho_{L}-(\mu_{W}+\mu_{B})-s}{\hat{\sigma}_{BL}^{(t)}+\sigma_{W}}\right)\leq\alpha$, we require
$\frac{\rho_{L}-(\mu_{W}+\mu_{B})-s}{\hat{\sigma}_{BL}^{(t)}+\sigma_{W}}\leq\Phi^{-1}(\alpha),$
@@
$s\geq\rho_{L}-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)\,(\hat{\sigma}_{BL}^{(t)}+\sigma_{W}).$
@@
$s_{t}^{*}=\max\Bigl{\{}0,\;\rho_{L}-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)\,(\hat{\sigma}_{BL}^{(t)}+\sigma_{W})\Bigr{\}}.$
@@
$\mathbb{P}\!\>\Bigl{(}\Pi_{j}^{(t)}+{s^{*}}^{(t)}<\rho_{L}\Bigr{)}\leq\alpha.$
Thus, the presence of ${s^{*}}^{(t)}$ makes lending feasible even under Bank $L$’s initial risk assessment $\hat{\sigma}_{BL}^{(t)}>\tilde{\sigma}_{L}$. Consequently, if the expected net profit (inclusive of the subsidy) is positive, Bank $L$ has an incentive to lend to group $B$.
Second, every time Bank $L$ extends a loan to group $B$, it observes a repayment drawn from $\mathcal{N}(\mu,\sigma_{B}^{2})$. Let $m$ be the number of such observations. Denote by $\hat{\sigma}_{B}^{2(m)}$ the sample variance computed from these $m$ observations. By the Strong Law of Large Numbers,
$\hat{\sigma}_{B}^{2(m)}\to\sigma_{B}^{2}\quad\text{almost surely as }m\to\infty.$
Because $\sigma_{B}<\tilde{\sigma}_{L}$, there exists a (finite) index $m^{*}$ (and thus a finite time $\tau$) such that for all $m\geq m^{*}$ the updated estimate satisfies:
$\sigma_{B}\leq{\hat{\sigma}_{B}^{(m)}}<\tilde{\sigma}_{L}.$
Third, once the updated variance estimate satisfies $\hat{\sigma}_{BL}^{(t)}<\tilde{\sigma}_{L}$, we can evaluate the required subsidy as
${s^{*}}^{(t)}=\max\Bigl{\{}0,\rho_{L}-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)\,(\hat{\sigma}_{BL}^{(t)}+\sigma_{W})\Bigr{\}}=0,$
by the same calculation as in Lemma 5. Thus, for all $t\geq\tau$, Bank $L$ is able to satisfy its $\mathrm{VaR}$ constraint without any subsidy. Since the number of observations required is finite almost surely, we conclude that with probability one there exists a finite $\tau$ such that for all $t\geq\tau$, ${s^{*}}^{(t)}=0$ and Bank $L$ continues lending to group $B$ without further external support.
@@
$\Pr(\Pi_{j}^{(t)}+G(t)<\rho_{L})\leq\alpha,$
where $G(t)$ is the guarantee provided in period $t$. After sufficient data collection, the updated variance estimate $\hat{\sigma}_{BL}^{(t)}$ will satisfy $\hat{\sigma}_{BL}^{(t)}<\tilde{\sigma}_{L}$, allowing Bank $L$ to lend without further guarantees.
@@
For any such $G(t)$, the $\mathrm{VaR}$ constraint is satisfied during each period $t$ of the exploration phase. Specifically, Bank $L$ is guaranteed that its effective return, $\pi_{B}^{(t)}+G(t)$, will not fall below $\rho_{L}$ with probability exceeding $\alpha$. This ensures that the bank has an incentive to lend to group $B$, provided the expected return is positive. The guarantee mechanism defined in the corollary generalizes the subsidy mechanism from Theorem 2. In the original subsidy framework, the guarantee function is given explicitly by $G(t)={s^{*}}^{(t)}=\max\{0,\rho_{L}-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)\,(\hat{\sigma}_{BL}^{(t)}+\sigma_{W})\}$, which satisfies the condition $\Pr(\Pi_{j}^{(t)}+{s^{*}}^{(t)}<\rho_{L})\leq\alpha$. The corollary allows for any $G(t)$ that satisfies the same probabilistic constraint. Once Bank $L$ updates its risk estimate to $\hat{\sigma}_{BL}^{(t)}<\tilde{\sigma}_{L}$, the $\mathrm{VaR}$ constraint is naturally satisfied without external support. Specifically, the guarantee $G(t)$ is no longer required, as the bank can safely lend to group $B$ on its own. Formally, this follows from Lemma 2.

### v2

@@ Appendix A Appendix
We derive the posterior distribution for $\sigma^{2}_{B}$ after observing $M$ loan outcomes.
Initial beliefs based on historical credit data.
Before observing any data, the bank starts with an uninformative prior:
$\sigma^{2}_{B}\sim\text{Inv-}\Gamma(a_{0},b_{0})$
The bank observes $np_{B}$ historical loan outcomes from group B applicants with complete files.
The likelihood of observing the historical data given $\sigma^{2}_{B}$ is:
$p(C_{B}|\sigma^{2}_{B})\propto(\sigma^{2}_{B})^{-np_{B}/2}\exp\left(-\frac{np_{B}}{2\sigma^{2}_{B}}\right)$
By Bayes’ theorem:
$p(\sigma^{2}_{B}|C_{B})\propto p(C_{B}|\sigma^{2}_{B})\cdot p(\sigma^{2}_{B})$
$\propto(\sigma^{2}_{B})^{-np_{B}/2}\exp\left(-\frac{np_{B}}{2\sigma^{2}_{B}}\right)\cdot(\sigma^{2}_{B})^{-a_{0}-1}\exp\left(-\frac{b_{0}}{\sigma^{2}_{B}}\right)$
$(\sigma^{2}_{B})^{-(a_{0}+\frac{np_{B}}{2})-1}\exp\left(-\frac{b_{0}+\frac{np_{B}}{2}}{\sigma^{2}_{B}}\right)$
Therefore, the initial belief is:
$\sigma^{2}_{B,0}\sim\text{Inv-}\Gamma\left(a_{0}+\frac{np_{B}}{2},b_{0}+\frac{np_{B}}{2}\right)$
Learning through lending.
Given the true variance $\sigma^{2}_{B}$ and known mean $\mu_{B}$, the observed returns $\{\pi_{B,1},\ldots,\pi_{B,M}\}$ are independently distributed:
$\pi_{B,m}|\sigma^{2}_{B}\sim\mathcal{N}(\mu_{B},\sigma^{2}_{B})\quad\text{for }m=1,\ldots,M$
The prior is:
$p(\sigma^{2}_{B})\propto(\sigma^{2}_{B})^{-{a_{0}+\frac{np_{B}}{2}}-1}\exp\left(-\frac{b_{0}+\frac{np_{B}}{2}}{\sigma^{2}_{B}}\right)$
The likelihood for $M$ observations with known mean $\mu_{B}$ is:
$p(\pi_{B,1},\ldots,\pi_{B,M}|\sigma^{2}_{B})=\prod_{m=1}^{M}\frac{1}{\sqrt{2\pi\sigma^{2}_{B}}}\exp\left(-\frac{(\pi_{B,m}-\mu_{B})^{2}}{2\sigma^{2}_{B}}\right)$
$(2\pi\sigma^{2}_{B})^{-M/2}\exp\left(-\frac{\sum_{m=1}^{M}(\pi_{B,m}-\mu_{B})^{2}}{2\sigma^{2}_{B}}\right)$
$\propto(\sigma^{2}_{B})^{-M/2}\exp\left(-\frac{\sum_{m=1}^{M}(\pi_{B,m}-\mu_{B})^{2}}{2\sigma^{2}_{B}}\right)$
By Bayes’ theorem:
$p(\sigma^{2}_{B}|\text{data})\propto p(\text{data}|\sigma^{2}_{B})\cdot p(\sigma^{2}_{B})$
$\propto(\sigma^{2}_{B})^{-M/2}\exp\left(-\frac{\sum_{m=1}^{M}(\pi_{B,m}-\mu_{B})^{2}}{2\sigma^{2}_{B}}\right)\cdot(\sigma^{2}_{B})^{-a_{0}-\frac{np_{B}}{2}-1}\exp\left(-\frac{b_{0}+\frac{np_{B}}{2}}{\sigma^{2}_{B}}\right)$
$(\sigma^{2}_{B})^{-(a_{0}+\frac{np_{B}+M}{2})-1}\exp\left(-\frac{b_{0}+\frac{np_{B}+\sum_{m=1}^{M}(\pi_{B,m}-\mu_{B})^{2}}{2}}{\sigma^{2}_{B}}\right)$
This is the kernel of an Inverse-Gamma distribution with parameters:
$a_{post}=a_{0}+\frac{np_{B}+M}{2},\quad b_{post}=b_{0}+\frac{np_{B}+\sum_{m=1}^{M}(\pi_{B,m}-\mu_{B})^{2}}{2}$
so that:
$\sigma^{2}_{B}|\pi_{B,1},\ldots,\pi_{B,M}\sim\text{Inv-}\Gamma\left(a_{0}+\frac{np_{B}+M}{2},b_{0}+\frac{np_{B}+\sum_{m=1}^{M}(\pi_{B,m}-\mu_{B})^{2}}{2}\right)$
The mean of an Inverse-Gamma(a,b)(a,b) distribution is $\frac{b}{a-1}$. Therefore:
$\mathbb{E}[\sigma^{2}_{B}|\pi_{B,1},\ldots,\pi_{B,M}]=\frac{b_{0}+\frac{np_{B}+\sum_{m=1}^{M}(\pi_{B,m}-\mu_{B})^{2}}{2}}{a_{0}+\frac{np_{B}+M}{2}-1}$
Which is the formula for the bank’s posterior estimate of the variance of lending to group $B$, given that it has observed M loans in addition to prior group-level credit history data. ∎
@@
First, we suppose that Bank $L$ lends to group $W$, and derive a threshold rule for the prior variance $\sigma^{2}_{B}$ holding $\sigma^{2}_{W}$ fixed such that this rule determines whether or not Bank $L$ will lend to group $B$.
@@
We then need to check Bank $L$’s VaRVaR constraint. It is:
$Pr\left(\Pi_{jt}<\rho\;|\;A_{BLt}=1\right)=\Phi\left(\dfrac{\rho_{L}-(\mu_{B}+\mu_{W})}{\hat{\sigma}_{Bt}+\sigma_{W}}\right)$
@@
$Pr\left(\Pi_{jt}\,|\,A_{BL}=1<\rho\right)\leq\alpha\>\Longleftrightarrow\Phi\left(\dfrac{\rho_{L}-(\mu_{B}+\mu_{W})}{\hat{\sigma}_{Bt}+\sigma_{W}}\right)\leq\alpha$
@@
$\hat{\sigma}^{2}_{Bt}\leq\left(\dfrac{\rho_{L}-(\mu_{W}+\mu_{B}+\Phi^{-1}(\alpha)\sigma_{W})}{\Phi^{-1}(\alpha)}\right)^{2}$
Where the right-hand side defines a variance threshold $\tilde{\sigma}^{2}_{L}$ such that, if group B’s prior variance is above this threshold, Bank $L$ will not lend to group $B$:
$\hat{\sigma}^{2}_{BLt}\leq\tilde{\sigma}^{2,\text{pool}}_{L}\equiv\left(\dfrac{\rho_{L}-(\mu_{W}+\mu_{B}+\Phi^{-1}(\alpha)\sigma_{W})}{\Phi^{-1}(\alpha)}\right)^{2}$
@@
$\sigma^{2}_{W}\leq\left(\dfrac{\rho_{L}-\mu_{W}}{\Phi^{-1}(\alpha)}\right)^{2}$
Since we know that, by Assumption 3, this condition is satisfied, Bank $L$ is willing to lend to $W$ unilaterally. Assumption 3 also tells us that $\hat{\sigma}^{2}_{BL0}$ is above this threshold, indicating that Bank $L$ will not lend to $B$ unilaterally.
Hence, Bank $L$ will lend to group $W$ unilaterally, and will lend to group $B$ in addition to group $W$ if and only if group $B$’s variance satisfies $\hat{\sigma}^{2}_{BLt}\leq\tilde{\sigma}^{2}_{L}$. ∎
Proof of Lemma 3
@@
$Pr\left(\Pi_{jt}<\rho|A_{BL}=1\right)=\Phi\left(\dfrac{\rho_{L}-[1+\nu_{t}](\mu_{B}+\mu_{W})}{\hat{\sigma}_{Bt}+\sigma_{W}}\right)$
@@
$\hat{\sigma}_{BHt}^{2}\leq\tilde{\sigma}^{2,\text{pool}}_{H}\equiv\left(\dfrac{\rho_{L}-[(1+\nu_{t})(\mu_{W}+\mu_{B})+\Phi^{-1}(\alpha)\sigma_{W}]}{\Phi^{-1}(\alpha)}\right)^{2}$
Proof of Corollary 1
@@
This can be deduced by inspection, but note that, by convention, $\alpha<.1$, so that $\Phi^{-1}(\alpha)<0$. ∎
@@
In each period, Bank L has belief, $\hat{\pi}_{Bt}\sim\mathcal{N}\bigl(\mu,\hat{\sigma}^{2}_{BLt}\bigr)$, so that:
$\mathbb{P}\!\left(\Pi_{jt}+s<\rho_{L}\right)=\Phi\!\left(\frac{\rho_{L}-(\mu_{W}+\mu_{B})-s}{\hat{\sigma}_{BLt}+\sigma_{W}}\right).$
Thus, to guarantee $\Phi\!\left(\frac{\rho_{L}-(\mu_{W}+\mu_{B})-s}{\hat{\sigma}^{2}_{BLt}+\sigma_{W}}\right)\leq\alpha$, we require
$\frac{\rho_{L}-(\mu_{W}+\mu_{B})-s}{\hat{\sigma}_{BLt}+\sigma_{W}}\leq\Phi^{-1}(\alpha),$
@@
$s\geq\rho_{L}-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)\,(\hat{\sigma}_{BLt}+\sigma_{W}).$
@@
$s_{t}^{*}=\max\Bigl\{0,\;\rho_{L}-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)\,(\hat{\sigma}_{BLt}+\sigma_{W})\Bigr\}.$
@@
$\mathbb{P}\!\>\Bigl(\Pi_{jt}+{s_{t}^{*}}<\rho\Bigr)\leq\alpha.$
Thus, the presence of $s_{t}^{*}$} makes lending feasible even under Bank $L$’s initial risk assessment $\hat{\sigma}^{2}_{BL0}>\tilde{\sigma}^{L}$. Consequently, if the expected net profit (inclusive of the subsidy) is positive, Bank $L$ has an incentive to lend to group $B$.
Second, every time Bank $L$ extends a loan to group $B$, it observes a repayment drawn from $\mathcal{N}(\mu,\sigma_{B}^{2})$. Let $m$ be the number of such observations. By the Strong Law of Large Numbers,
$\hat{\sigma}_{BLm}^{2}\to\sigma_{B}^{2}\quad\text{almost surely as }m\to\infty.$
Because $\sigma^{2}_{B}<\tilde{\sigma}^{2,\text{pool}}_{L}$, there exists a (finite) index $m^{*}$ (and thus a finite time $\tau$) such that for all $m\geq m^{*}$ the updated estimate satisfies:
$\sigma^{2}_{B}\leq\hat{\sigma}^{2}_{BLm}<\tilde{\sigma}^{2,\text{pool}}_{L}.$
Third, once the updated variance estimate satisfies $\hat{\sigma}^{2}_{BLt}<\tilde{\sigma}^{2,\text{pool}}_{L}$, we can evaluate the required subsidy as
${s^{*}}^{(t)}=\max\Bigl\{0,\rho_{L}-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)\,(\hat{\sigma}_{BLt}+\sigma_{W})\Bigr\}=0,$
by the same calculation as in Lemma 5. Thus, for all $t\geq\tau$, Bank $L$ is able to satisfy its $\mathrm{VaR}$ constraint without any subsidy. Since the number of observations required is finite almost surely, we conclude that with probability one there exists a finite $\tau$ such that for all $t\geq\tau$, ${s_{t}^{*}}=0$ and Bank $L$ continues lending to group $B$ without further external support.
@@
$\Pr(\Pi_{jt}+G(t)<\rho)\leq\alpha,$
where $G(t)$ is the guarantee provided in period $t$. After sufficient data collection, the updated variance estimate $\hat{\sigma}^{2}_{BLt}$ will satisfy $\hat{\sigma}^{2}_{BLt}<\tilde{\sigma}^{2,\text{pool}}_{L}$, allowing Bank $L$ to lend without further guarantees.
@@
For any such $G(t)$, the $\mathrm{VaR}$ constraint is satisfied during each period $t$ of the exploration phase. Specifically, Bank $L$ is guaranteed that its effective return, $\pi_{B}^{(t)}+G(t)$, will not fall below $\rho$ with probability exceeding $\alpha$. This ensures that the bank has an incentive to lend to group $B$, provided the expected return is positive. The guarantee mechanism defined in the corollary generalizes the subsidy mechanism from Theorem 2. In the original subsidy framework, the guarantee function is given explicitly by $G(t)={s_{t}^{*}}=\max\{0,\rho_{L}-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)\,(\hat{\sigma}_{BLt}+\sigma_{W})\}$, which satisfies the condition $\Pr(\hat{\Pi}_{jt}+{s^{*}_{t}}<\rho)\leq\alpha$. The corollary allows for any $G(t)$ that satisfies the same probabilistic constraint. Once Bank $L$ updates its risk estimate to $\hat{\sigma}^{2}_{BLt}<\tilde{\sigma}^{2,\text{pool}}_{L}$, the $\mathrm{VaR}$ constraint is naturally satisfied without external support. Specifically, the guarantee $G(t)$ is no longer required, as the bank can safely lend to group $B$ on its own. Formally, this follows from Lemma 2.
@@
Extension to Expected Shortfall
Variance Thresholds under Expected Shortfall
We now derive the variance threshold under Expected Shortfall and show it is more conservative than the VaR constraint. The logic of the argument is the same throughout, however.
When Bank L lends to both groups W and B:
$\Pi_{L}=\pi_{W}+\pi_{B}$
where from the bank’s perspective (using its estimate $\hat{\sigma}^{2}_{B}$):
$\Pi_{L}\sim\mathcal{N}(\mu_{W}+\mu_{B},\sigma^{2}_{W}+\hat{\sigma}^{2}_{B})$
$\text{VaR}_{\phi}(\Pi_{L})=(\mu_{W}+\mu_{B})+\sigma_{W}+\hat{\sigma}_{BLt}\cdot\Phi^{-1}(\phi)$
Using the definition $\text{ES}_{\alpha}(X)=\frac{1}{\alpha}\int_{0}^{\alpha}\text{VaR}_{\phi}(X)d\phi$:
$\text{ES}_{\alpha}(\Pi_{L})$ $=\frac{1}{\alpha}\int_{0}^{\alpha}\left[(\mu_{W}+\mu_{B})+\sigma_{W}+\hat{\sigma}_{BLt}\cdot\Phi^{-1}(\phi)\right]d\phi$
$=(\mu_{W}+\mu_{B})+\frac{\sigma_{W}+\hat{\sigma}_{BLt}}{\alpha}\int_{0}^{\alpha}\Phi^{-1}(\phi)d\phi$
Using the fact that $\int_{0}^{\alpha}\Phi^{-1}(\phi)d\phi=-\phi(\Phi^{-1}(\alpha))$:
$\text{ES}_{\alpha}(\Pi_{L})=(\mu_{W}+\mu_{B})-\sigma_{W}+\hat{\sigma}_{BLt}\cdot\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}$
Bank L requires $\text{ES}_{\alpha}(\Pi_{L})\geq\rho_{L}$:
$(\mu_{W}+\mu_{B})-\sigma_{W}+\hat{\sigma}_{BLt}\cdot\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}\geq\rho_{L}$
Rearranging for $\hat{\sigma}^{2}_{B}$:
$\sigma_{W}+\hat{\sigma}_{BLt}\leq\frac{(\mu_{W}+\mu_{B})-\rho_{L}}{\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}}$
Squaring both sides:
$\sigma^{2}_{W}+\hat{\sigma}^{2}_{B}\leq\left(\frac{(\mu_{W}+\mu_{B})-\rho_{L}}{\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}}\right)^{2}$
Therefore:
$\hat{\sigma}^{2}_{B}\leq\left(\frac{(\mu_{W}+\mu_{B})-\rho_{L}}{\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}}\right)^{2}-\sigma^{2}_{W}\equiv\tilde{\sigma}^{2,\text{pool},ES}_{L}$
From Lemma 3, the VaR threshold is:
$\tilde{\sigma}^{2,\text{pool},\text{VaR}}_{L}=\left(\frac{\rho_{L}-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)\sigma_{W}}{\Phi^{-1}(\alpha)}\right)^{2}$
We next show that $\tilde{\sigma}^{2,\text{pool},ES}_{L}<\tilde{\sigma}^{2,\text{pool},\text{VaR}}_{L}$, and hence, that the ES threshold is more restrictive.
Since $\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}>\Phi^{-1}(\alpha)$, the denominator in the ES threshold calculation is larger, yielding:
$\tilde{\sigma}^{2,\text{pool},ES}_{L}<\tilde{\sigma}^{2,\text{pool},\text{VaR}}_{L}$
This shows that Expected Shortfall creates a more conservative variance threshold than VaR, making it harder for Bank L to lend to group B initially.
We can solve for the optimal subsidy under Expected Shortfall analogously:
Optimal Subsidy under Expected Shortfall
In order to lend to group B, Bank L requires:
$\text{ES}_{\alpha}(\Pi_{L}+s)\geq\rho$
where Expected Shortfall with subsidy $s$ is:
$\text{ES}_{\alpha}(\Pi_{Lt}+s)=(\mu_{W}+\mu_{B}+s)-(\sigma_{W}+\hat{\sigma}_{BLt})\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}$
Rearranging for $s$:
$s\geq\rho-(\mu_{W}+\mu_{B})+(\sigma_{W}+\hat{\sigma}_{BLt})\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}$
The optimal (minimum) subsidy is:
$s^{*}_{ES,t}=\max\left\{0,\rho-(\mu_{W}+\mu_{B})+(\sigma_{W}+\hat{\sigma}_{BLt})\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}\right\}$
The VaR-derived subsidy from Lemma 6 is:
$s^{*}_{VaR,t}=\max\left\{0,\rho-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)(\sigma_{W}+\hat{\sigma}_{BLt})\right\}$
Since $\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}>\Phi^{-1}(\alpha)$ for $\alpha<.1$, (e.g. 2.063 ¿ 1.645 when $\alpha=0.05$):
$s^{*}_{ES,t}>s^{*}_{VaR,t}$
The Expected Shortfall-based subsidy is larger, reflecting the fact that Expected Shortfall is a more conservative risk criterion than Value-at-Risk.